Carson Wood Products processes logs into four grades of lumber totaling 500,000 board feet as follows at a joint cost of $300,000:

Grade Board Feet Final Sales Value
First and second 75,000 $ 56,250
No. 1 common 200,000 180,000
No. 2 common 100,000 105,000
No. 3 common 125,000 127,500


What amount of joint costs will be allocated to No. 2 common using the constant gross margin percentage method?

a.$192,000

b.$67,200

c.$37,800

d.$300,000
Answer #1

Gross margin ratio = Gross margin / revenue

Revenue = 56250+180000+105000+127500= 468,750

Gross margin = 468,750-300,000 = 168,750

Gross margin ratio = 168,750/468,750 = 36%

Joint costs allocated to no 2 = 105,00 * 36% = 37,800

Option C is the answer